OpenStudy (anonymous):

How do I solve this binomial (factoring) : x^8-256x^4. Thank you :)

OpenStudy (mertsj):

First factor out the common factor which is x^4

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

x^4 *(x^2 - 16) *(x^2 + 16) = x^4 *(x-4)(x+4)(x^2 + 16)

OpenStudy (anonymous):

Is this the answer: x^4(x^2+16)?

OpenStudy (mertsj):

\[x ^{4}(x ^{4}-16) = x ^{4}(x ^{2}-4)(x ^{2}+4)=x ^{4}(x-2)(x+2)(x ^{2}+4)\]

OpenStudy (anonymous):

So its the last one?

OpenStudy (mertsj):

x4(x−2)(x+2)(x2+4)

OpenStudy (anonymous):

Is a formula used to solve this?

OpenStudy (mertsj):

The difference of two squares.

OpenStudy (anonymous):

how did you get (x^2+4)

OpenStudy (mertsj):

Use it twice. Once on the x^4-16 and again on x^2-4

OpenStudy (mertsj):

\[x ^{4}-16=(x ^{2}-4)(x ^{2}+4)\]

OpenStudy (anonymous):

@Mertsj .. it should be (x-4)(x+4) instead of (x-2)(x+2) .. you have to factor 256 which is 16^2

OpenStudy (anonymous):

x^4 - 256 = (x - 4)(x+4)(x^2 + 16)

OpenStudy (anonymous):

dont you take x^2 out of x^4 Mertsj

OpenStudy (mertsj):

Add a factor of x^2+16)

OpenStudy (mertsj):

\[x ^{4}-256=(x ^{2}-16)(x ^{2}+16)=\]

OpenStudy (anonymous):

so wouldnt the final answer not have x^4, what is the forumula exactly?

OpenStudy (mertsj):

\[(x-4)(x+4)(x ^{2}+16)\]

OpenStudy (mertsj):

Then put the x^4 that we factored out in the front of it.

OpenStudy (mertsj):

\[x ^{4}(x-4)(x+4)(x ^{2}+16)\]

OpenStudy (mertsj):

That's the final answer.

OpenStudy (anonymous):

cant you break down the 16 tho?

OpenStudy (mertsj):

No. The sum of two squares is not factorable in the set of real numbers.

OpenStudy (mertsj):

The difference of two squares is factorable but not the sum.

OpenStudy (anonymous):

Thanks want to help me with another? :)

OpenStudy (anonymous):

its set up different

OpenStudy (mertsj):

Sure

OpenStudy (anonymous):

9x^2+64

OpenStudy (mertsj):

That is the sum of two squares and it cannot be factored.

OpenStudy (anonymous):

So its prime?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

ok, be right back :3

OpenStudy (anonymous):

im going to start a new question, this time i want to check an answer i got :)